Vincent is 10 years older than John Paul.Five years ago,his age was twice that of John Paul.Find their present age.
Answer by palanisamy(496)

Let the age of Vincent = x.
Then the age of John Paul = x-10
Five years ago,Vincent's age was twice that of John Paul.
x-5 = 2(x-15)
x-5 = 2x-30
x-2x= -30+5
-x = -25
x = 25
At present, the age of Vincent = 25.
And the age of John Paul = x-10 = 25-10 = 15